OpenStudy (anonymous):

MEDALS Functions f(x) and g(x) are shown below. f(x) = x2 g(x) = x2 - 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)?

OpenStudy (anonymous):

Options Left by 4 units Right by 4 units Left by 8 units Right by 8 units

OpenStudy (anonymous):

g(x) = x^2 - 8x + 16 You can see that the above function is a perfect square. Hope you remember that (a-b)^2=a^2-2ab+b^2 x^2 - 8x + 16 is the same with a=x and b=4 x^2 - 8x + 16=(x-4)^2 So the shift from x^2 is by 4 units. And it is to the right. You can see that easily of you sketch the graphs.

OpenStudy (anonymous):

thanks so much

OpenStudy (secret-ninja):

""So the shift from x^2 is by 4 units. And it is to the right. You can see that easily of you sketch the graphs. I put B and I got it wrong :o

OpenStudy (lalaland_lauren):

It's to the left by four units, because the four is negative and it shouldn't be going in the positive direction.